40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog, Page 225 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  6675   Thu May 24 14:49:59 2012 KojiSummaryGeneralDaily news idea

Top tab categolies:

  • Summary
  • CDS
    • CDS Status
  • PEM
    • Seismic 24h trend
    • Accoustic 24h trend
    • Weather/Temp/Barometer/etc 24h trend
  • PSL/IOO
    • PSL summary trend / duty ratio
    • IOO summary (MC Health Check/IOO QPD trends / IFO QPD trends / Transmon QPD trends) duty ratio
  • SUS
    • Summary
    • OSEM PSD/trend
    • OPLEV PSD/trend
  • IFO
    • DC Mon
    • RF ports
    • OMC
  • Steve
    • Vacuum
  • Misc.

IFO

  • DC Monitors
    • Incident beam power trend (24h)
    • AS/REFL/POP/TRX/TRY bean power trend (24h)
    • AS/POP RF beam power trend (24h)
  • RF port
    • DARM sensitivity PSD (mean/min/max/reference) for an hour
    • DARM/CARM/PRCL/MICH/SRCL PSD
    • DARM/CARM/PRCL/MICH/SRCL (freq vs Gaussianity)
    • DARM/CARM/PRCL/MICH/SRCL calibration trend
  • OMC
    • TBD

 

  6674   Thu May 24 13:28:38 2012 DenUpdatePEMHEPA

HEPA filter was running at 90% of max. I reduced it to 20%. Acoustic noise moved down

psl_acoustic.png

The range of MCL oscillations has also decreased but fluctuations in the frequency range 10-100 are still present.

MCL is much more stable now.

mcl_90.png        mcl_20.png

 

  6673   Thu May 24 12:23:00 2012 DenUpdateGeneralconnection error monitor for the website

The main page of connection error monitor can be a scheme of computers and models with connections. If there is an error in connection, the corresponding arrows turns from green to red.

errors_overview.png

 

Each arrow is a line to a more detailed information about the channels.

errors.png

  6672   Thu May 24 10:10:04 2012 steveUpdateIOOPMC locked

 The PMC  behavior is not changed.

 

Attachment 1: pmc4d.png
pmc4d.png
  6671   Thu May 24 02:55:36 2012 DenUpdatePEMacoustic noise in MCL

Mic in the PSL showed that fluctuations in the MCL in the frequency range 10 - 100 Hz are due to acoustic noise. I've measured MCL, MCL / PSL mic coherence 2 times with interval 300 seconds.

Surprisingly, acoustic noise level did not change but MC sometimes is more sensitive to acoustic noise, sometimes less.

audio_noise.png

  6670   Thu May 24 01:17:13 2012 DenUpdateCDSPMC autolocker

Quote:

 

  • SCRIPT
    • Auto-locker for PMC, PSL things - DEN

 I wrote auto-locker for PMC. It is called autolocker_pmc, located in the scripts directory, svn commited. I connected it to the channel C1:PSL-PMC_LOCK.  It is currently running on rosalba. MC autolocker runs on op340m, but I could not execute the script on that machine

op340m:scripts>./autolock_pmc
./autolock_pmc: Stale NFS file handle.

I did several tests, usually, the script locks PMC is a few seconds. However, if PMC DC output has been drift significantly, if might take longer as discussed below.

The algorithm:

       if autolocker if enabled, monitor PSL-PMC_PMCTRANSPD channel
       if TRANS is less then 0.4, start locking:
               disengage PMC servo by enabling PMC TEST 1
               change PSL-PMC_RAMP unless TRANS is higher then 0.4 (*)
               engage PMC servo by disabling PMC TEST 1
       else sleep for 1 sec
 

(*) is tricky. If RAMP (DC offset) is specified then TRANS will be oscillating in the range ( TRANS_MIN, TRANS_MAX ). We are interested only in the TRANS_MAX. To make sure, we estimate it right, TRANS channel is read 10 times and the maximum value is chosen. This works good.

Next problem is to find the proper range and step to vary DC offset RAMP. Of coarse, we can choose the maximum range (-7, 0) and minimum step 0.0001, but it will take too long to find the proper DC offset. For that reason autolocker tries to find a resonance close to the previous DC offset in the range (RAMP_OLD - delta, RAMP_OLD + delta), initial delta is 0.03 and step is 0.003. It resonance is not found in this region, then delta is multiplied by a factor of 2 and so on. During this process RAMP range is controlled not to be wider then (-7, 0).

The might be a better way to do this. For example, use the gradient descent algorithm and control the step adaptively. I'll do that if this realization will be too slow.

I've disabled autolocker_pmc for the night.

  6669   Wed May 23 21:32:15 2012 SureshUpdateIOOWFS didn't turn off automatically

Quote:

I just sat down in the control room, and discovered the PMC (and everything else) unlocked.  I relocked the PMC, but the MC wasn't coming back.  After a moment of looking around, I discovered that the WFS were on, and railing.  I ran the "turn WFS off" script, and the MC came back right away, and the WFS came on as they should.

We need to relook at the WFS script, or the MC down script, to make sure that any time the MC is unlocked, no matter why it unlocked, the WFS output is off and the filter histories are cleared.

    The only script that can currently take this action is the MC autolocker.  If that is disabled first and the PMC unlocks later, the WFS will not be turned off.  During the last round of discussions we had about the autolocker script, sometime last Nov, we decided that too much automation is not desirable and that the autolocker should be kept as simple as possible.

 

  6668   Wed May 23 20:41:37 2012 SureshBureaucracyGeneral40m Meeting Action Items: Tip-tilts : cabling and electronics

Quote:
..........
  • Tip Tilts
    • Prepare electronics for TTs (coil drivers) - JAMIE
    • In-air TT testing to confirm we can control / move TTs before we vent - SURESH
    • Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH
  • .....

 

 [Koji, Suresh]

  We tried to locate the sixteen analog output channels we need to control the four tip-tilts (four coils on each).  We have only 8 available channels on the C1SUS machine. 

  So we will have to plug-in a new DAC output card on one of the machines and it would be logical to do that on the C1IOO machine as the active tip-tilts are conceptually part of the IOO sub-system.   We have to procure this card if we do not already have it.  We have to make an interface between this card output and a front panel on the 1X2 rack.  We may have to move some of the sub-racks on the 1X2 rack to accommodate this front panel.

  We checked out the availability of cards (De-whitening, Anti-imaging, SOS coil drivers)  yesterday.  In summary: we have all the cards we need (and some spares too).  As the De-whitening and Anti-imaging cards each have 8 channels, we need only two of each to address the sixteen channels.  And we need four of the SOS coil drivers, one for each tip-tilt.  There are 9 slots available on the C1IOO satellite expansion chassis (1X1 rack), where these eight cards could be accommodated.

  There are two 25 pin feed-thoughs, where the PZT drive signals currently enter the BS chamber.  We will have to route the SOS coil driver outputs to these two feed-throughs. 

   Inside the BS chamber, there are cables which carry the PZT signals from the chamber wall to the the table top, where they are anchored to a post (L- bracket).  We need a 25-pin-to-25-pin cable (~2m length) to go from the post to the tip-tilt (one for each tip-tilt).  And then, of course, we need quadrapus cables (requested from Rich) which fit inside each tip-tilt to go to the BOSEMs.

   I am summarising it all here to give an overview of the work involved.

 

 

  6667   Wed May 23 13:57:11 2012 JenneBureaucracyGeneral40m Meeting Action Items

PRIORITY ITEMS:

Align Ygreen beam - JENNE, YUTA

Arm cavity sweeps, mode scan - JENNE, YUTA

ASS doesn't run on Ubuntu! or CentOS Fix it! - YUTA, JENNE, JAMIE's help

Input matricies, output filters to tune SUS.  check after upgrade. - JENNE

POX11 whitening is not toggling the analog whitening??? - JAMIE, JENNE, KOJI

Decide on plots for 40m Summary page - DEN, STEVE, JENNE, KOJI, JAMIE, YUTA, SURESH, RANA, DUNCAN from Cardiff/AEI

 Look into PMC PZT drift - PZT failing? Real MC length change? - JENNE, KOJI, YUTA

THE FULL LIST:

  • General
    • Revamp control room - more monitors, maybe move TV shelves. - JENNE
    • Audio system for the signals!!!! Even a crappy one! - LEO
  • SUS
    • Input matricies, output filters to tune SUS.  check after upgrade. - JENNE
  • IFO
    • Fix occasional common-mode power transient in the arm transmissions. Probably an alignment thing. Would ISS help? - MIKEJ?
    • Drift of the green incident axis -> Assess the amount of the drift / replace the mount - JENNE, KOJI
    • Calibration of POP22 / AS110 - JENNE
    • PMC/IMC/ARM characterization (loss, finesse, reflectivity, etc) - JENNE
    • Align green beam - JENNE, YUTA
    • Arm cavity sweeps, mode scan - JENNE, YUTA
    • Align AS OSA (others?) - JENNE  
    • Investigate PRMI glitches, instability - JENNE
    • PZT or Picomotor mounts for PSL/ALS beams - WHO??
    • ALS on the both arm simultaneously / common / diff ALS scripts - JENNE
  • CDS
    • Capture OSA signals in CDS (the 'scope TDS1001B has a USB port in the back for connecting to the computer) - WHO??
    • Transmon (arms) for high and low power - WHO??
    • POX11 whitening is not toggling the analog whitening??? - JAMIE, JENNE, KOJI
    • Install guardians to monitor EPICS values - WHO??
  • Electronics
    • Actuator noise level characterization (coil driver response in V/m & coil driver noise level V/rtHz) - WHO??
    • Beat box - KOJI
    • I/Q DFD - WHO??
    • Improvement of POP22/110/AS110 RF circuits. - WHO??
  • MEDM
    • Complete 40m overview screen - everything should be clickable with pseudo 3D icons - JENNE
    • Better suspension screen - copy from sites. - WHO??
    • Script to generate a MEDM tree - WHO??
    • Resurrect MEDM snapshots - WHO??
    • New ! buttons on every screen, include wiki page - WHO??
    • Screen for errors in signal transmission between models - DEN
  • SCRIPT
    • Locking scripts with integrator triggers / disabling when unlocked - JENNE
    • Daily diagnosis of the MC spot positions (there must be something already...) - SURESH?
    • Daily/occasional adjustment of the incident axis on the MC - SURESH?
    • IFO_CONFIGURE scripts still do a burt restore, so step the optics.  Need to remove optic alignment from config .snap files, use reg restore script instead. - JENNE
    • OPLEV/OSEM trending script before the IFO work for diagnosis. Put into 40m summary screen. - JENNE, DUNCAN from Cardiff/AEI
    • Auto-locker for arms - JENNE
    • Auto-locker for PMC, PSL things - DEN
    • Diagnostic script for CDS - mx_stream, other stuff. - WHO??
  • Video
    • If each video screen has a caption, that would be great - WHO??
    • GUI interface of "videoswitch" - MIKEJ
    • Mouse Jiggler for Zita (called Swinput?) - JENNE
  • Ubuntu vs. CentOS
    • burttoday is not aliased in ubuntu.    burttoday:      aliased to cd /opt/rtcds/caltech/c1/burt/autoburt/today/ - JAMIE
    • Upgrade Allegra, make sure connect to DTT, Dataviewer, AWG. - JAMIE
    • ASS doesn't run on Ubuntu! or CentOS Fix it! - JENNE, JAMIE's help
    • CentOS machines cannot open simulink models properly (get "Bad Link"s everywhere, so you can't do anything useful) - JAMIE
  • MM
    • IPPOS beam measurement - SURESH with JENNE
    • AS beam measurement (if beam is bright enough) - SURESH and JENNE
    • Mode matching calculations, sensitivity to MC waist measurement errors, PRM position - JENNE
    • Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT - JAMIE, JENNE, KOJI, SURESH
    • Use sensoray to capture, measure beam mode at AS, POP - YUTA
  • Stray Light
    • Scattered light measurement at the end stations: design / confirmation of the mechanical parts/optics/cameras - JAN
    • Align AUX laser into dark port - KOJI, YUTA
  • OAF
    • OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN
    • Static-only OAF noise budget (Adaptive noise budget as next step) - DEN
    • Script for daily / weekly re-calculation of Wiener, post to elog if need changing - DEN
  • Tip Tilts
    • Prepare electronics for TTs (coil drivers) - JAMIE
    • In-air TT testing to confirm we can control / move TTs before we vent - SURESH
    • Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH
  • RF Photodiodes
    • Opto Energy diode laser - purchase - ERICG
    • Purchase fibers, splitters, other supplies - ERICG
    • Set everything up - ERICG

 

  6666   Wed May 23 11:30:39 2012 steveUpdatePEMnew emergency back up lights installed

New self checking emergency back up lights were installed at 13 locations in the 40m lab.

Attachment 1: IMG_1217.JPG
IMG_1217.JPG
  6665   Wed May 23 10:40:21 2012 steveUpdateIOOPMC locked

Quote:

I locked the PMC and the MC followed instantly.

 

Attachment 1: pmc.png
pmc.png
  6664   Tue May 22 23:36:02 2012 DenUpdatePEMmircophones

I've put 5 EM172 microphones close together and measured there signal and coherence. They are plugged in to accelerometer channels.

close_micro.png

 

Then I've suspended microphones around the MC - 2 at MC2, 2 at MC1,3 and 1 at PSL. The amplifier box is above STS readout box.

 DSC_4288.JPG           DSC_4299.JPG       DSC_4304.JPG      DSC_4302.JPG

Microphone close to PSL gave a strong coherence with MC_F, as we already saw it using Blue Bird Microphone.

ACC_MC2_XY channels <=> MC2 microphones

ACC_MC1_XY channels <=> MC1,3 microphones

ACC_MC1_Z channel <=> PSL microphones

5_mic.png

 

  6663   Tue May 22 20:46:38 2012 yutaUpdateCDSMEDM suspension screens using macro

I fixed the problem Jamie pointed out in elog #6657 and #6659.

What I did:
1. Created the following template files in /opt/rtcds/userapps/trunk/sus/c1/medm/templates/ directry.

SUS_SINGLE_LOCKIN1.adl
SUS_SINGLE_LOCKIN2.adl
SUS_SINGLE_LOCKIN_INMTRX.adl
SUS_SINGLE_OPTLEV_SERVO.adl
SUS_SINGLE_PITCH.adl
SUS_SINGLE_POSITION.adl
SUS_SINGLE_SUSSIDE.adl
SUS_SINGLE_TO_COIL_MASTER.adl
SUS_SINGLE_COIL.adl
SUS_SINGLE_YAW.adl
SUS_SINGLE_INMATRIX_MASTER.adl
SUS_SINGLE_INPUT.adl
SUS_SINGLE_TO_COIL_X_X.adl
SUS_SINGLE_OPTLEV_IN.adl
SUS_SINGLE_OLMATRIX_MASTER.adl

To open these files, you have to define $(OPTIC) and $(DCU_ID).
For SUS_SINGLE_TO_COIL_X_X.adl, you also have to define $(FILTER_NUMBER), too. See SUS_SINGLE_TO_COIL_MASTER.adl.

2. Fixed the following screens so that they open SUS_SINGLE.adl.

C1SUS_WATCHDOGS.adl
C1IOO_MC_ALIGN.adl
C1IOO_WFS_MASTER.adl
C1IFO_ALIGN.adl

  6662   Tue May 22 20:24:06 2012 JamieUpdateComputersrossa is now running Ubuntu 10.04

Now same as pianosa and rosalba.  I'll upgrade allegra on Friday.

  6661   Tue May 22 20:01:26 2012 DenUpdateCDSerror monitor

I've created transmission error monitors in rfm, oaf, sus, lsc, scx, scy and ioo models. I tried to get data from every channel transmitted through PCIE and RFM. I also included some shared memory channels.

The medm screen is in the EF STATUS -> TEM. It shows 16384 for the channels that come from simulation plant. Others are 0, that's fine.

  6660   Tue May 22 13:06:11 2012 JenneUpdateIOOWFS didn't turn off automatically

I just sat down in the control room, and discovered the PMC (and everything else) unlocked.  I relocked the PMC, but the MC wasn't coming back.  After a moment of looking around, I discovered that the WFS were on, and railing.  I ran the "turn WFS off" script, and the MC came back right away, and the WFS came on as they should.

We need to relook at the WFS script, or the MC down script, to make sure that any time the MC is unlocked, no matter why it unlocked, the WFS output is off and the filter histories are cleared.

  6659   Tue May 22 11:47:43 2012 JamieUpdateCDSMEDM suspension screens using macro

Actually, it looks like we're not quite done here.  All the paths in the SUS_SINGLE screen need to be updated to reflect the move.  We should probably make a macro that points to /opt/rtcds/caltech/c1/screens, and update all the paths accordingly.

  6658   Tue May 22 11:45:12 2012 JamieConfigurationCDSPlease remember to commit SVN changes

Hey, folks.  Please remember to commit all changes to the SVN in a timely manor.  If you don't, multiple commits will get lumped together and we won't have a good log of the changes we're making.  You might also end up just loosing all of your work.  SVN COMMIT when you're done!  But please don't commit broken or untested code.

pianosa:release 0> svn status | grep -v '^?'
M       cds/c1/models/c1rfm.mdl
M       sus/c1/models/c1mcs.mdl
M       sus/c1/models/c1scx.mdl
M       sus/c1/models/c1scy.mdl
M       isc/c1/models/c1lsc.mdl
M       isc/c1/models/c1pem.mdl
M       isc/c1/models/c1ioo.mdl
M       isc/c1/models/ADAPT_XFCODE_MCL.c
M       isc/c1/models/c1oaf.mdl
M       isc/c1/models/c1gcv.mdl
M       isc/common/medm/OAF_OVERVIEW.adl
M       isc/common/medm/OAF_DOF_BLRMS.adl
M       isc/common/medm/OAF_OVERVIEW_BAK.adl
M       isc/common/medm/OAF_ADAPTATION_MICH.adl
pianosa:release 0>

  6657   Tue May 22 11:32:02 2012 JamieUpdateCDSMEDM suspension screens using macro

Very nice, Yuta!  Don't forget to commit your changes to the SVN.  I took the liberty of doing that for you.  I also tweaked the file a bit, so we don't have to specify IFO and SYS, since those aren't going to ever change.  So the arguments are now only: OPTIC=MC1,DCU_ID=36.  I updated the sitemap accordingly.

Yuta, if you could go ahead and modify the calls to these screens in other places that would be great.  The WATCHDOG, LSC_OVERVIEW, MC_ALIGN screens are ones that immediately come to mind.

And also feel free to make cool new ones.  We could try to make simplified version of the suspension screens now being used at the sites, which are quite nice.

  6656   Tue May 22 08:03:01 2012 steveUpdateSUSETMY oplev gets new laser

Quote:

 

 The typical sign of a dying gas laser is that it glows for a few minutes only. The power supplies are fine.

Two new  JDS - Uniphase 1103P lasers ( NT64-104 )  arriving on Monday, May 21

 Yesterday I swapped in new He/Ne laser with output power 3.5 mW  The return spot on qpd is large ~6mm in diameter and 20,500 counts

The spot size reduction require similar layout as ETMX oplev.

  6655   Tue May 22 00:23:45 2012 DenUpdateCDStransmission error monitor

I've started to create channels and an medm screen to monitor the errors that occur during the transmission through the RFM model. The screen will show the amount of lost data per second for each channel.

Not all channels are ready yet. For created channels, number of errors is 0, this is good.

 Screenshot.png

  6654   Mon May 21 21:27:39 2012 yutaUpdateCDSMEDM suspension screens using macro

Background:
 We need more organized MEDM screens. Let's use macro.

What I did:
1. Edited /opt/rtcds/userapps/trunk/sus/c1/medm/templates/SUS_SINGLE.adl using replacements below;

sed -i s/#IFO#SUS_#PART_NAME#/'$(IFO)$(SYS)_$(OPTIC)'/g SUS_SINGLE.adl
sed -i s/#IFO#SUS#_#PART_NAME#/'$(IFO)$(SYS)_$(OPTIC)'/g SUS_SINGLE.adl
sed -i s/#IFO#:FEC-#DCU_ID#/'$(IFO):FEC-$(DCU_ID)'/g SUS_SINGLE.adl
sed -i s/#CHANNEL#/'$(IFO):$(SYS)-$(OPTIC)'/g SUS_SINGLE.adl
sed -i s/#PART_NAME#/'$(OPTIC)'/g SUS_SINGLE.adl

2. Edited sitemap.adl so that they open SUS_SINGLE.adl with arguments like
    IFO=C1,SYS=SUS,OPTIC=MC1,DCU_ID=36
instead of opening ./c1mcs/C1SUS_MC1.adl.

3. I also fixed white blocks in the LOCKIN part.

Result:
  Now you don't have to generate every suspension screens. Just edit SUS_SIGNLE.adl.

Things to do:
 - fix every other MEDM screens which open suspension screens, so that they open SUS_SINGLE.adl
 - make SUS_SINGLE.adl more cool

  6653   Mon May 21 19:16:55 2012 KojiUpdateGeneralrossa X config

I rebooted rossa and the X sever has stalled.

  6652   Mon May 21 07:45:39 2012 steveUpdateIOOPMC locked

I locked the PMC and the MC followed instantly.

Attachment 1: pmclocked.png
pmclocked.png
  6651   Sun May 20 19:57:51 2012 DenUpdatePEMmicrophones

I've soldered EM172 microphones to BNC connectors to get data from them.

DSC_4287.JPG    DSC_4285.JPG

Then I've build an amplifier for them. The circuit is

scheme.png

I've build 6 such circuits inside 1 box. It needs +15 V to A3 and GND to A2. A1 power channel is not used.

DSC_4284.JPG     DSC_4286.JPG

LISO model for this scheme was created and simulation results were compared to measurement of each channel

freq_resp.png   noise.png

Measured noise curve (green) is the SR785 own noise.

  6650   Fri May 18 15:25:15 2012 steveUpdateSUSAR coated lens swapped in at ETMX oplev

Quote:

 ETMX oplev had 6 mm diameter beam on the qpd.  I relayed the beam path with 2 lenses  to get back  3 mm beam on the qpd

BRC 037  -100 Bi _concave lens and PCX 25  200 VIS do the job. Unfortunately the concave lens has the AR 1064.

 

 

 The uncoated bi-concave lens was replaced by AR coated one: KBC 037 -100 AR.14 resulting 35% count increase on qpd

Attachment 1: ETMXoplevlensAR.png
ETMXoplevlensAR.png
  6649   Fri May 18 11:19:17 2012 steveUpdateSUSETMY oplev laser is dead

 

 The typical sign of a dying gas laser is that it glows for a few minutes only. The power supplies are fine.

Two new  JDS - Uniphase 1103P lasers ( NT64-104 )  arriving on Monday, May 21

  6648   Thu May 17 15:31:22 2012 steveUpdateGeneralbeam trap posts & clamps for vacuum

 Aluminum posts and SS clamps for green glass traps in vacuum will be out of the shops by June 6, 2012 the latest

 

Attachment 1: 05161201.PDF
05161201.PDF 05161201.PDF
Attachment 2: 05151201.PDF
05151201.PDF
  6647   Wed May 16 14:06:49 2012 steveUpdateSUSTemperature rules

Guralp 1  is on the south side of IOO chamber

Attachment 1: temprules.png
temprules.png
  6646   Wed May 16 11:53:45 2012 JenneUpdateComputersImage Subtraction

Quote:

It doesn't look all that different, but the first image didn't have that much lit up in it to begin with.

 This is totally cool!  You can see that the OSEM lights are almost entirely gone in the subtracted image.

Can you switch to trying with one of the *TM*F cameras?  (ITMXF, ITMYF, ETMYF, ETMXF)  They tend to have more background, so there should be a more dramatic subtraction.  Den or Suresh should be able to lock one of the arms for you.

  6645   Tue May 15 23:40:46 2012 Mike J.UpdateComputersImage Subtraction

I acquired 2 raw frames of MC2 using "/users/mjenson/sensoray/sdk_2253_1.2.2_linux/capture -n -s 720x480 -f 1", one while the laser was off the mode cleaner and another while it was on:

mc2_1.bmp mc2_2.bmp

I then used "/users/mjenson/sensoray/sdk_2253_1.2.2_linux/imsub/display-image.py" to generate bitmaps of the raw images, which I then subtracted using the Python Imaging Library to generate a new image:

mc2_1-mc2_2.bmp

It doesn't look all that different, but the first image didn't have that much lit up in it to begin with. I should be able to write a script that does all of this without needing to generate new files in between acquisition and subtraction.

  6644   Tue May 15 08:44:55 2012 steveUpdateSUSETMX damping restored

ETMX sus damping restored

Attachment 1: ETMX_ETMYvar.png
ETMX_ETMYvar.png
  6643   Mon May 14 08:49:47 2012 steveUpdateSUSETMY damping restored

ETMY sus damping restored.

  6642   Fri May 11 23:33:41 2012 DenUpdateAdaptive Filteringoffline vs online

I've compared offline Wiener filtering with online static + adaptive filtering for MC_F with GUR1_XYZ and GUR2_XYZ as witness signals

off_on.jpg

Note: online filter works up to 32 Hz (AI filter at 32 Hz is used). There is no subtraction up from this frequency, just MC_F was measured in different times for online and offline filtering. This difference in MC_F in frequency range 20-100 Hz showed up again as before with microphone testing.  One can see it in 1 minute. Smth is noisy.

Reasons why online filter is worse then offline:

1. FIR -> IIR conversion produces error. Now I'm using VECTFIT approximation with 16 poles (splitting into 2 filter banks), this not enough. I tried to use 50 and split them into 5 filter banks, but this scheme is not working: zpk -> sos conversion produces error and the result filter works completely wrong.

2. Actuator TF. VECTFIT works very good here - we have only 1 resonance. However, it should be measured precisely.

3. Account for AA and AI filters that rotate the phase at 1-10 Hz by ~ 10 degrees.

  6641   Fri May 11 17:21:56 2012 JenneUpdateIOOOAF left enabled - MC unlocked for more than an hour

No leaving the OAF running until you're sure (sure-sure, not kind of sure, not pretty sure, but I've enabled guardians to make sure nothing bad can happen, and I've been sitting here watching it for 24+ hours and it is fine) that it works okay.

OAF (both adaptive and static paths) were left enabled, which was kicking MC2 a lot.  Not quite enough that the watchdog tripped, but close.  The LSCPOS output for MC2 was in the 10's or 100's of thousands of counts.  Not okay.

This brings up the point though, which I hadn't actively thought through before, that we need an OAF watchdog.  An OAF ogre?  But a benevolent ogre.  If the OAF gets out of control, it's output should be shut off.  Eventually we can make it more sophisticated, so that it resets the adaptive filter, and lets things start over, or something. 

But until we have a reliable OAF ogre, no leaving the adaptive path enabled if you're not sitting in the control room.  The static path should be fine, since it can't get out of control on it's own.

Especially no leaving things like this enabled without a "I'm leaving this enabled, I'll be back in a few hours to check on it" elog! 

  6640   Fri May 11 08:07:30 2012 JamieUpdateCDSFB

Quote:

Already for the second time today all computers loose connection to the framebuilder. When I ssh to framebuilder DAQD process was not running. I started it

controls@fb ~ 130$ sudo /sbin/init q

Just to be clear, "init q" does not start the framebuilder.  It just tells the init process to reparse the /etc/inittab.  And since init is supposed to be configured to restart daqd when it dies, it restarted it after the reloading of /etc/inittab.  You and Alex must have forgot to do that after you modified the inittab when you're were trying to fix daqd last week.

daqd is known to crash without reason.  It usually just goes unnoticed because init always restarts it automatically.  But we've known about this problem for a while.

Quote:

But I do not know what causes this problem. May be this is a memory issue. For FB

Mem:   7678472k total,  7598368k used,    80104k free

Practically all memory is used. If more is needed and swap is off, DAQD process may die.

This doesn't really mean anything, since the computer always ends up using all available memory.  It doesn't indicate a lack of memory.  If the machine is really running out of memory you would see lots of ugly messages in dmesg.

  6639   Thu May 10 22:05:21 2012 DenUpdateCDSFB

Already for the second time today all computers loose connection to the framebuilder. When I ssh to framebuilder DAQD process was not running. I started it

controls@fb ~ 130$ sudo /sbin/init q

But I do not know what causes this problem. May be this is a memory issue. For FB

Mem:   7678472k total,  7598368k used,    80104k free

Practically all memory is used. If more is needed and swap is off, DAQD process may die.

  6638   Thu May 10 21:13:01 2012 DenUpdateElectronicsADC 3

ADC 3 INPUT 4 (#3 in the c1pem model if you count from 0) is bad. It adds DC = ~1 V to the signal as well as noise. I plugged in GUR2 channels to STS1 channels (7-9).

  6637   Thu May 10 14:49:06 2012 KojiHowToGeneralHow to clean & bake black glass pieces

  6636   Wed May 9 15:50:13 2012 JenneBureaucracyGeneral40m Meeting Action Items

I'm combining the IFO check-up list (elog 6595) and last week's action items list (elog 6597).  I thought about making it a wiki page, but this way everyone has to at least scroll past the list ~1/week.

Feel free to cross things out as you complete them, but don't delete them.  Also, if there's a WHO?? and you feel inspired, just do it!

PRIORITY ITEMS:

Dither-align arm to get IR on actuation nodes, align green beam - JENNE

Arm cavity sweeps, mode scan - JENNE

ASS doesn't run on Ubuntu! or CentOS Fix it! - JENNE, JAMIE's help

Input matricies, output filters to tune SUS.  check after upgrade. - JENNE

POX11 whitening is not toggling the analog whitening??? - JAMIE, JENNE, KOJI

OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN

 

THE FULL LIST:

  • General
    • Revamp control room - more monitors, maybe move TV shelves. - JENNE
    • Audio system for the signals!!!! Even a crappy one! - LEO
  • SUS
    • ETMX oplev - replace 1064nm lens with 632nm lens (KBC037, -100) - JENNE
    • Input matricies, output filters to tune SUS.  check after upgrade. - JENNE
  • IFO
    • Fix occasional common-mode power transient in the arm transmissions. Probably an alignment thing. Would ISS help? - WHO??
    • Drift of the green incident axis -> Assess the amount of the drift / replace the mount - JENNE, KOJI
    • Calibration of POP22 / AS110 - JENNE
    • PMC/IMC/ARM characterization (loss, finesse, reflectivity, etc) - JENNE
    • Dither-align arm to get IR on actuation nodes, align green beam - JENNE
    • Arm cavity sweeps, mode scan - JENNE
    • Align AS OSA (others?) - JENNE  
    • Investigate PRMI glitches, instability - JENNE
    • PZT or Picomotor mounts for PSL/ALS beams - WHO??
    • ALS on the both arm simultaneously / common / diff ALS scripts - JENNE
  • CDS
    • Capture OSA signals in CDS (the 'scope TDS1001B has a USB port in the back for connecting to the computer) - WHO??
    • Transmon (arms) for high and low power - WHO??
    • POX11 whitening is not toggling the analog whitening??? - JAMIE, JENNE, KOJI
    • Install guardians to monitor EPICS values - WHO??
  • Electronics
    • Actuator noise level characterization (coil driver response in V/m & coil driver noise level V/rtHz) - WHO??
    • Beat box - KOJI
    • I/Q DFD - WHO??
    • Improvement of POP22/110/AS110 RF circuits. - WHO??
  • MEDM
    • Complete 40m overview screen - everything should be clickable with pseudo 3D icons - JENNE
    • Better suspension screen - copy from sites. - WHO??
    • Script to generate a MEDM tree - WHO??
    • Resurrect MEDM snapshots - WHO??
    • New ! buttons on every screen, include wiki page - WHO??
  • SCRIPT
    • Locking scripts with integrator triggers / disabling when unlocked - JENNE
    • Daily diagnosis of the MC spot positions (there must be something already...) - SURESH?
    • Daily/occasional adjustment of the incident axis on the MC - SURESH?
    • IFO_CONFIGURE scripts still do a burt restore, so step the optics.  Need to remove optic alignment from config .snap files, use reg restore script instead. - JENNE
    • OPLEV/OSEM trending script before the IFO work for diagnosis. - JENNE
    • Auto-locker for arms - JENNE
    • Auto-locker for PMC, PSL things - DEN
    • Diagnostic script for CDS - mx_stream, other stuff. - WHO??
  • Video
    • If each video screen has a caption, that would be great - WHO??
    • GUI interface of "videoswitch" - MIKEJ
    • Mouse Jiggler for Zita (called Swinput?) - JENNE
  • Ubuntu vs. CentOS
    • burttoday is not aliased in ubuntu.    burttoday:      aliased to cd /opt/rtcds/caltech/c1/burt/autoburt/today/ - JAMIE
    • Upgrade Rossa, Allegra, make sure they connect to DTT, Dataviewer, AWG. - JAMIE
    • ASS doesn't run on Ubuntu! or CentOS Fix it! - JENNE, JAMIE's help
    • CentOS machines cannot open simulink models properly (get "Bad Link"s everywhere, so you can't do anything useful) - JAMIE
  • MM
    • IPPOS beam measurement - SURESH with JENNE
    • AS beam measurement (if beam is bright enough) - SURESH and JENNE
    • Mode matching calculations, sensitivity to MC waist measurement errors, PRM position - JENNE
    • Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT - JAMIE, JENNE, KOJI, SURESH
    • Use sensoray to capture, measure beam mode at AS, POP - YUTA
  • Stray Light
    • Scattered light measurement at the end stations: design / confirmation of the mechanical parts/optics/cameras - JAN
  • OAF
    • OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN
    • Static-only OAF noise budget (Adaptive noise budget as next step) - DEN
    • Build amplifiers for new small microphones - DEN
    • Script for daily / weekly re-calculation of Wiener, post to elog if need changing - DEN
  • Tip Tilts
    • Prepare electronics for TTs (coil drivers) - JAMIE
    • In-air TT testing to confirm we can control / move TTs before we vent - SURESH
    • Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH
  • RF Photodiodes
    • Opto Energy diode laser - look into buying - ERICG
    • Purchase fibers, splitters, other supplies - ERICG
    • Set everything up - ERICG

  6635   Wed May 9 15:02:50 2012 DenUpdateCDSRFM

Quote:

However, this did not help, C1RFM did not start. I decided to restart all models on C1SUS machine in hope that C1RFM uses some other models and can't connect to them but this suspended C1SUS machine.

 This happened because of the code bug -

// If PCIE comms show errors, may want to add this cache flushing
#if 1
if(ipcInfo[ii].netType == IPCIE)
          clflush_cache_range (&(ipcInfo[ii].pIpcData->dBlock[sendBlock][ipcIndex].data), 16); // & was missing - Alex fixed this
#endif
 

After this bug was fixed and the code was recompiled, C1:OAF_MCL_IN is OK, no errors occur during the transmission C1:OAF-MCL_ERR=0.

So the problem was in the PCIE card that could not send such amount of data and the last channel (MCL is the last) was corrupted. Now, when Alex added cache flushing, the problem is fixed.

We should spend some more attention to such problems. This time 2046 out of 2048 points were lost per second. But what if 10-20 points are lost, we would not notice that in the dataviewer, but this will cause problems.

  6634   Wed May 9 14:32:31 2012 JenneUpdateCDSBurt restored

Den and Alex left things not-burt restored, and Den mentioned to me that it might need doing.

I burt restored all of our epics.snaps to the 1am today snapshot.  We lost a few hours of striptool trends on the projector, but now they're back (things like the BLRMS don't work if the filters aren't engaged on the PEM model, so it makes sense).

  6633   Wed May 9 11:31:50 2012 DenUpdateCDSRFM

I added PCIE memory cache flushing to c1rfm model by changing 0 to 1 in /opt/rtcds/rtscore/release/src/fe/commData2.c on line 159, recompiled and restarted c1rfm.

Jamie, do not be mad at me, Alex told me do that!

However, this did not help, C1RFM did not start. I decided to restart all models on C1SUS machine in hope that C1RFM uses some other models and can't connect to them but this suspended C1SUS machine. After reboot encounted the same C1SUS -> FB communication error and fixed it in the same was as in the previous case of C1SUS reboot. This happens already the second time (out of 2) after C1SUS machine reboot.

I changed /opt/rtcds/rtscore/release/src/fe/commData2.c back, recompiled and restarted c1rfm. Now everything is back. C1RFM -> C1OAF is still bad.

  6632   Wed May 9 10:46:54 2012 DenUpdateCDSOAF model not seeing MCL correctly

Quote:

Den noticed this, and will write more later, I just wanted to sum up what Alex said / did while he was here a few minutes ago....

 From my point of view during rfm -> oaf transmission through Dolphin we loose a significant part of the signal. To check that I've created MEDM screen to monitor the transmission errors in the OAF model. It shows how many errors occurs per second. For MCL channel this number turned out to be 2046 +/- 1. This makes sense to me as the sampling rate is 2048 Hz => then we actually receive only 1-3 data points per second. We can see this in the dataviewer.

C1:OAF-MCL_IN follows C1:IOO-MC_F in the sense that the scale of 2 signals are the same in 2 states: MC locked and unlocked. It seems that we loose 2046 out of 2048 points per second.

oaf_rec.png

  6631   Wed May 9 09:19:22 2012 KojiUpdateLockingPRM is really moving when PRMI locked

Is this enhancement of spectrum caused by the lock? Or by the actuation?

If this is also seen with approximately same amount of actuation to PRM POS,
this is just a suspension problem.

If this is only seen with the PRM locked, this is somehow related to the opt-mechanical coupling.

  6630   Wed May 9 08:21:42 2012 JamieUpdateCDSNo signals for DTT from SUS

Quote:

 c1iscey is much less happy - neither the IOP nor the scy model are willing to talk to fb.  I might give up on them after another few minutes, and wait for some daytime support, since I wanted to do DRMI stuff tonight.

Yeah, giving up now on c1iscey (Jamie....ideas are welcome).  I can lock just fine, including the Yarm, I just can't save data or see data about ETMY specifically.  But I can see LSC data, so I can lock, and I can now take spectra of corner optics.

 This is the mx_stream issue reported previously.  The symptom is that all models on a single front end loose contact with the frame builder, as opposed to *all* models on all front end loosing contact with the frame builder.  That indicates that the problem is a common fb communication issue on the single front end, and that's all handled with mx_stream.

ssh'ing into c1iscey and running "sudo /etc/init.d/mx_stream restart" fixed the problem.

  6629   Wed May 9 04:47:20 2012 JenneUpdateLockingPRM is really moving when PRMI locked

A few things tonight.  Locked both arms simultaneously (IR only).  Locked MICH, Locked PRMI, although it doesn't like staying locked for more than a minute or so, and not always that long.

Locking PRCL was possible by getting rid of the power normalization.  We need to get some triggering going on for the power norm.  I think it's a good idea for after the cavity is locked, but when PRCL is not locked, POP22 is ~0, so Refl33/Pop22 is ~inf.  The PRCL loop kept railing at the Limit that was set.  Getting rid of the power normalization fixed this railing. 

I took some spectra of PRM's oplev, while PRMI was locked, and unlocked.  The PRM is definitely moving more when the cavity is locked.  I'm not sure yet what to do about this, but the result was repeatable many times (~6 or 7 over an hour or so).  The OpLev spectra when PRMI was locked didn't depend too strongly on the PRM's alignment, although I think that's partly because I wasn't able to really get the PRM to optimal alignment.  I think POP22I is supposed to get to 7 or so...last week with Koji it was at least flashing that high.  But tonight I couldn't get POP22I above 4, and most of the time it wouldn't go above 3.  As I was aligning PRM and the circulating SB power increased, POP22I fluctuations increased significantly, then the cavity unlocked.  So maybe this is because as I get closer, PRM gets more wiggly.  I tried playing 'chicken' with it, and took spectra as I was aligning PRM (align, get some improvement, stop to take spectra, then align more, stop to take spectra....) but usually it would fall out of lock after 1-2 iterations of this incremental alignment and I'd have to start over.  When it relocked, it usually wouldn't come back to the same level of POP22I, which was kind of disappointing. 

In the PDF attached, pink and light blue are when the PRMI is locked, and red and dark blue are no PRCL feedback.  The effect is more pronounced with Pitch, but it's there for both Pitch and Yaw.

Also, I need to relook at my new restore/misalign scripts.  They were acting funny tonight, so I'm taking back my "they're awesome, use them without thinking about it" certification.

Attachment 1: PRM_louder_when_aligned.pdf
PRM_louder_when_aligned.pdf
  6628   Wed May 9 01:14:44 2012 JenneUpdateCDSNo signals for DTT from SUS

Quote:

UPDATE UPDATE:  Genius me just checked the FE status screen again.  It was fine ~an hour ago when I sat down to start interferometer-izing for the night, but now the SUS model and both of the ETMY computer models are having problems connecting to the fb.  *sigh* 

 Restarted SUS model - it's now happy. 

c1iscey is much less happy - neither the IOP nor the scy model are willing to talk to fb.  I might give up on them after another few minutes, and wait for some daytime support, since I wanted to do DRMI stuff tonight.

Yeah, giving up now on c1iscey (Jamie....ideas are welcome).  I can lock just fine, including the Yarm, I just can't save data or see data about ETMY specifically.  But I can see LSC data, so I can lock, and I can now take spectra of corner optics.

  6627   Wed May 9 00:45:13 2012 JenneUpdateCDSNo signals for DTT from SUS

Upgrades suck.  Or at least making everything work again after the upgrade.

On the to-do list tonight:  look at OSEM sensor and OpLev spectra for PRM, when PRMI is locked and unlocked.  Goal is to see if the PRM is really moving wildly ("crazy" as Kiwamu always described it) when it's nicely aligned and PRMI is locked, or if it's an artifact of lever arm between PRM and the cameras (REFL and AS).

However, I can't get signals on DTT.  So far I've checked a bunch of signals for SUS-PRM, and they all either (a) are just digital 0 or (b) are ADC noise.  Lame.

Steve's elog 5630 shows what reasonable OpLev spectra should look like:  exactly what you'd expect.

Attached below is a small sampling of different SUS-PRM signals.  I'm going to check some other optics, other models on c1sus, etc, to see if I can narrow down where the problem is.  LSC signals are fine (I checked AS55Q, for example).

UPDATE:  SRM channels are same ADC noise.  MC1 channels are totally fine.  And Den had been looking at channels on the RFM model earlier today, which were fine.

ETMY channels - C1:SUS-ETMY_LLCOIL_IN1 and C1:SUS-ETMY_SUSPOS_IN1 both returned "unable to obtain measurement data".  OSEM sensor channels and OpLev _PERROR channel were digital zeros.

ETMX channels were fine

UPDATE UPDATE:  Genius me just checked the FE status screen again.  It was fine ~an hour ago when I sat down to start interferometer-izing for the night, but now the SUS model and both of the ETMY computer models are having problems connecting to the fb.  *sigh* 

Attachment 1: OpLevs_OSEMs_allSUSchannels_ADCnoiseOnly.pdf
OpLevs_OSEMs_allSUSchannels_ADCnoiseOnly.pdf
Attachment 2: Screenshot.png
Screenshot.png
  6626   Tue May 8 17:48:50 2012 JenneUpdateCDSOAF model not seeing MCL correctly

Den noticed this, and will write more later, I just wanted to sum up what Alex said / did while he was here a few minutes ago....

 

Errors are probably really happening.... c1oaf computer status 4-bit thing:  GRGG.  The Red bit is indicating receiving errors.  Probably the oaf model is doing a sample-and-hold thing, sampling every time (~1 or 2 times per sec) it gets a successful receive, and then holding that value until it gets another successful receive. 

Den is adding EPICS channels to record the ERR out of the PCIE dolphin memory CDS_PART, so that we can see what the error is, not just that one happened.

Alex restarted oaf model:  sudo rmmod c1oaf.ko, sudo insmod c1oaf.ko .  Clicked "diag reset" on oaf cds screen several times, nothing changed.  Restarted c1oaf again, same rmmod, insmod commands.

Den, Alex and I went into the IFO room, and looked at the LSC computer, SUS computer, SUS I/O chassis, LSC I/O chassis and the dolphin switch that is on the back of the rack, behind the SUS IO chassis.  All were blinking happily, none showed symptoms of errors.

Alex restarted the IOP process:  sudo rmmod c1x04, sudo insmod c1x04.  Chans on dataviewer still bad, so this didn't help, i.e. it wasn't just a synchronization problem.  oaf status: RRGG. lsc status: RGGG. ass status: RGGG.

sudo insmod c1lsc.ko, sudo insmod c1ass.ko, sudo insmod c1oaf.ko .  oaf status: GRGG. lsc status: GGGG. ass status: GGGG.  This means probably lsc needs to send something to oaf, so that works now that lsc is restarted, although oaf still not receiving happily.

Alex left to go talk to Rolf again, because he's still confused.

Comment, while writing elog later:  c1rfm status is RRRG, c1sus status is RRGG, c1oaf status is GRGG, both c1scy and c1scx are RGRG.  All others are GGGG.

ELOG V3.1.3-